OpenStudy (anonymous):

4.A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

solve two equations \[x-0.525=.0.1\] and \[x-0.525=-0.01\]

OpenStudy (anonymous):

both are solved by adding \(0.0525\) to both sides

OpenStudy (kash_thesmartguy):

Solve the equations:\[x+(-0.525+0.525)=-0.1+0.525\]and\[x+(-0.525+0.525)=-0.01+0.525\]Doing that will result in:\[x=0.425\]and\[x=0.515\]So then: x=0.425 and x=0.515